simtl2

0.1

Generated by Doxygen 1.5.5

Fri May 30 23:06:20 2008

Contents

1 Directory Hierarchy 1 1.1 Directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Namespace Index 3 2.1 Namespace List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3 Class Index 5 3.1 Class Hierarchy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

4 Class Index 7 4.1 Class List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

5 File Index 9 5.1 File List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

6 Directory Documentation 11 6.1 Desktop/ Directory Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 6.2 Desktop/Study/Programming/KDEProject/ Directory Reference . . . . . . . . . . 12 6.3 Desktop/Study/Programming/ Directory Reference . . . . . . . . . . . . . . . . . . 13 6.4 Desktop/Study/Programming/KDEProject/simtl2/ Directory Reference . . . . . . 14 6.5 Desktop/Study/Programming/KDEProject/simtl2/src/ Directory Reference …. 15 6.6 Desktop/Study/ Directory Reference . . . . . . . . . . . . . . . . . . . . . . . . . . 16

7 Namespace Documentation 17 7.1 Ui Namespace Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

8 Class Documentation 19 8.1 Analyzer Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 8.2 AnalyzerInterface Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 8.3 Arp_analysis Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 8.4 DataLink_analysis Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 8.5 DB_control Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 8.6 Icmp_analysis Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 8.7 Ip_analysis Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 8.8 Packet_analysis Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 8.9 Pcap_option Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 8.10 Simtl_dir Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 8.11 Simtl_gui Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 8.12 Tcp_analysis Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 8.13 Time_control Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 8.14 Ui::MainWindow Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 8.15 Ui_MainWindow Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

9 File Documentation 53

9.1 Desktop/Study/Programming/KDEProject/simtl2/src/analyzer_interface.h File

Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 9.2 Desktop/Study/Programming/KDEProject/simtl2/src/main.cpp File Reference . . 54 9.3 Desktop/Study/Programming/KDEProject/simtl2/src/moc_simtl_dir.cpp File

Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 9.4 Desktop/Study/Programming/KDEProject/simtl2/src/moc_simtl_gui.cpp File

Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 9.5 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.cpp File

Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 9.6 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.h File Ref-

erence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 9.7 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_db.cpp File Reference 60 9.8 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_db.h File Reference . 61 9.9 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_dir.cpp File Reference 62 9.10 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_dir.h File Reference 63 9.11 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_gui.cpp File Reference 64 9.12 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_gui.h File Reference 65 9.13 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_time.cpp File Refer-

ence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 9.14 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_time.h File Reference 67 9.15 Desktop/Study/Programming/KDEProject/simtl2/src/ui_simtl_gui.h File Refer-

ence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

Chapter 1

Directory Hierarchy

1.1 Directories

This directory hierarchy is sorted roughly, but not completely, alphabetically:

Desktop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

Study . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Programming . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 KDEProject . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 simtl2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 src . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

2  Directory Hierarchy

Chapter 2

Namespace Index

2.1 Namespace List

Here is a list of all namespaces with brief descriptions: Ui . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

4  Namespace Index

Class Index

3.1 Class Hierarchy

This inheritance list is sorted roughly, but not completely, alphabetically:

Analyzer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 AnalyzerInterface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 DB_control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Packet_analysis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

Arp_analysis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 DataLink_analysis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Icmp_analysis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Ip_analysis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 Tcp_analysis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

Pcap_option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 Simtl_dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 Simtl_gui . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Time_control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 Ui_MainWindow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

Ui::MainWindow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

6  Class Index

Class Index

4.1 Class List

Here are the classes, structs, unions and interfaces with brief descriptions:

Analyzer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 AnalyzerInterface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Arp_analysis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 DataLink_analysis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 DB_control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Icmp_analysis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Ip_analysis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 Packet_analysis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Pcap_option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 Simtl_dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 Simtl_gui . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Tcp_analysis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 Time_control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 Ui::MainWindow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 Ui_MainWindow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

8  Class Index

File Index

5.1 File List

Here is a list of all files with brief descriptions:

Desktop/Study/Programming/KDEProject/simtl2/src/analyzer_interface.h …. 53 Desktop/Study/Programming/KDEProject/simtl2/src/main.cpp . . . . . . . . . . . . 54 Desktop/Study/Programming/KDEProject/simtl2/src/moc_simtl_dir.cpp . . . . . 55 Desktop/Study/Programming/KDEProject/simtl2/src/moc_simtl_gui.cpp . . . . . 56 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.cpp . . . . . . 57 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.h . . . . . . . 58 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_db.cpp . . . . . . . . . 60 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_db.h . . . . . . . . . . 61 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_dir.cpp . . . . . . . . . 62 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_dir.h . . . . . . . . . . 63 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_gui.cpp . . . . . . . . . 64 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_gui.h . . . . . . . . . . 65 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_time.cpp . . . . . . . . 66 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_time.h . . . . . . . . . 67 Desktop/Study/Programming/KDEProject/simtl2/src/ui_simtl_gui.h . . . . . . . . 68

10  File Index

Chapter 6

Directory Documentation

6.1 Desktop/ Directory Reference

Directories

• directory Study

6.2 Desktop/Study/Programming/KDEProject/ Directory Reference

Directories

• directory simtl2

6.3 Desktop/Study/Programming/ Directory Reference  13

6.3 Desktop/Study/Programming/ Directory Reference

Directories

• directory KDEProject

6.4 Desktop/Study/Programming/KDEProject/simtl2/ Directory Reference

Directories

• directory src

6.5 Desktop/Study/Programming/KDEProject/simtl2/src/ Directory Reference 15

6.5 Desktop/Study/Programming/KDEProject/simtl2/src/ Directory Reference

Files

• file analyzer_interface.h

• file main.cpp

• file moc_simtl_dir.cpp

• file moc_simtl_gui.cpp

• file simtl_analysis.cpp

• file simtl_analysis.h

• file simtl_db.cpp

• file simtl_db.h

• file simtl_dir.cpp

• file simtl_dir.h

• file simtl_gui.cpp

• file simtl_gui.h

• file simtl_time.cpp

• file simtl_time.h

• file ui_simtl_gui.h

16  Directory Documentation

6.6 Desktop/Study/ Directory Reference

Directories

• directory Programming

Chapter 7

Namespace Documentation

7.1 Ui Namespace Reference

Classes

• class MainWindow

18  Namespace Documentation

Class Documentation

8.1 Analyzer Class Reference

#include <simtl_analysis.h>

Static Public Member Functions

• static void ether_Analysis (u_char ∗user, const struct pcap_pkthdr ∗h, const u_char ∗p)

8.1.1 Member Function Documentation

8.1.1.1 void Analyzer::ether_Analysis (u_char ∗ user, const struct pcap_pkthdr ∗ h, const u_char ∗ p) [static]

References Packet_analysis::data_analysis(),  db_control,  Pcap_option::get_chcnt(),

DataLink_analysis::get_ether_type(), DB_control::get_file(), Ip_analysis::get_ip_hlen(), Ip_analysis::get_ip_proto(), Ip_analysis::get_ip_totlen(), DB_control::get_table(), Pcap_- option::plus_chcnt(), DB_control::row_plus(), and DB_control::set_model().

Referenced by init_Analysis().

The documentation for this class was generated from the following files:

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.h

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.cpp

20  Class Documentation

8.2 AnalyzerInterface Class Reference

#include <analyzer_interface.h>

Public Member Functions

• AnalyzerInterface ()

• virtual ∼AnalyzerInterface ()

Private Member Functions

• AnalyzerInterface (const AnalyzerInterface &source)

• void operator= (const AnalyzerInterface &source)

8.2.1 Constructor & Destructor Documentation

8.2.1.1 AnalyzerInterface::AnalyzerInterface () [inline] 8.2.1.2 virtual AnalyzerInterface::∼AnalyzerInterface () [inline, virtual] 8.2.1.3 AnalyzerInterface::AnalyzerInterface (const AnalyzerInterface & source)

        [private]

8.2.2 Member Function Documentation

8.2.2.1 void AnalyzerInterface::operator= (const AnalyzerInterface & source)

        [private]

The documentation for this class was generated from the following file:

• Desktop/Study/Programming/KDEProject/simtl2/src/analyzer_interface.h

8.3 Arp_analysis Class Reference

#include <simtl_analysis.h>

Inheritance diagram for Arp_analysis::

Public Member Functions

• Arp_analysis (struct ether_header ∗ep)

• unsigned char get_arp_hwl (void)

• unsigned short get_arp_hwt (void)

• unsigned short get_arp_op (void)

• unsigned char get_arp_prl (void)

• unsigned short get_arp_prt (void)

• char ∗ get_arp_send_haddr (void)

• char ∗ get_arp_send_paddr (void)

• char ∗ get_arp_target_haddr (void)

• char ∗ get_arp_target_paddr (void)

Protected Attributes

• unsigned char arp_hwl

• unsigned short arp_hwt

• unsigned short arp_op

• unsigned char arp_prl

• unsigned short arp_prt

• char ∗ arp_send_haddr

• char ∗ arp_send_paddr

• char ∗ arp_target_haddr

• char ∗ arp_target_paddr

Private Attributes

• struct arphdr ∗ arph

8.3.1 Constructor & Destructor Documentation

8.3.1.1 Arp_analysis::Arp_analysis (struct ether_header ∗ ep)

References arp_hwl, arp_hwt, arp_op, arp_prl, arp_prt, arp_send_haddr, arp_send_paddr, arp_target_haddr, arp_target_paddr, arph, DB_control::column_plus(), db_control, Pcap_- option::get_chcnt(), DB_control::get_column(), DB_control::get_file(), DB_control::get_- itemModel(), DB_control::get_rows(), DB_control::get_sum(), Time_control::get_time(), Packet_analysis::p_point, and time_control.

22  Class Documentation

8.3.2 Member Function Documentation

8.3.2.1 unsigned char Arp_analysis::get_arp_hwl (void) [inline] References arp_hwl.

8.3.2.2 unsigned short Arp_analysis::get_arp_hwt (void) [inline] References arp_hwt.

8.3.2.3 unsigned short Arp_analysis::get_arp_op (void) [inline] References arp_op.

8.3.2.4 unsigned char Arp_analysis::get_arp_prl (void) [inline] References arp_prl.

8.3.2.5 unsigned short Arp_analysis::get_arp_prt (void) [inline] References arp_prt.

8.3.2.6 char∗ Arp_analysis::get_arp_send_haddr (void) [inline] References arp_send_haddr.

8.3.2.7 char∗ Arp_analysis::get_arp_send_paddr (void) [inline] References arp_send_paddr.

8.3.2.8 char∗ Arp_analysis::get_arp_target_haddr (void) [inline] References arp_target_haddr.

8.3.2.9 char∗ Arp_analysis::get_arp_target_paddr (void) [inline] References arp_target_paddr.

8.3.3 Member Data Documentation

8.3.3.1 unsigned char Arp_analysis::arp_hwl [protected] Referenced by Arp_analysis(), and get_arp_hwl().

8.3.3.2 unsigned short Arp_analysis::arp_hwt [protected] Referenced by Arp_analysis(), and get_arp_hwt().

Generated on Fri May 30 23:06:20 2008 for simtl2 by Doxygen

8.3 Arp_analysis Class Reference  23

8.3.3.3 unsigned short Arp_analysis::arp_op [protected] Referenced by Arp_analysis(), and get_arp_op().

8.3.3.4 unsigned char Arp_analysis::arp_prl [protected] Referenced by Arp_analysis(), and get_arp_prl().

8.3.3.5 unsigned short Arp_analysis::arp_prt [protected] Referenced by Arp_analysis(), and get_arp_prt().

8.3.3.6 char∗ Arp_analysis::arp_send_haddr [protected] Referenced by Arp_analysis(), and get_arp_send_haddr().

8.3.3.7 char∗ Arp_analysis::arp_send_paddr [protected] Referenced by Arp_analysis(), and get_arp_send_paddr().

8.3.3.8 char∗ Arp_analysis::arp_target_haddr [protected] Referenced by Arp_analysis(), and get_arp_target_haddr().

8.3.3.9 char∗ Arp_analysis::arp_target_paddr [protected] Referenced by Arp_analysis(), and get_arp_target_paddr().

8.3.3.10 struct arphdr∗ Arp_analysis::arph [read, private] Referenced by Arp_analysis(). The documentation for this class was generated from the following files:

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.h

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.cpp

24  Class Documentation

8.4 DataLink_analysis Class Reference

#include <simtl_analysis.h>

Inheritance diagram for DataLink_analysis::

Public Member Functions

• DataLink_analysis (struct ether_header ∗ep)

• char ∗ get_ether_dest ()

• char ∗ get_ether_source ()

• unsigned short get_ether_type ()

Protected Attributes

• char eth_destination [ETH_ALEN]

• char eth_source [ETH_ALEN]

• unsigned short ether_type

8.4.1 Constructor & Destructor Documentation

8.4.1.1 DataLink_analysis::DataLink_analysis (struct ether_header ∗ ep) References db_control, eth_destination, eth_source, ether_type, Pcap_option::get_chcnt(), DB_control::get_file(), DB_control::get_sum(), and Packet_analysis::i.

8.4.2 Member Function Documentation

8.4.2.1 char∗ DataLink_analysis::get_ether_dest () [inline] References eth_destination.

8.4.2.2 char∗ DataLink_analysis::get_ether_source () [inline] References eth_source.

8.4.2.3 unsigned short DataLink_analysis::get_ether_type () [inline] References ether_type. Referenced by Analyzer::ether_Analysis().

Generated on Fri May 30 23:06:20 2008 for simtl2 by Doxygen

8.4 DataLink_analysis Class Reference  25

8.4.3 Member Data Documentation

8.4.3.1 char DataLink_analysis::eth_destination[ETH_ALEN] [protected] Referenced by DataLink_analysis(), and get_ether_dest().

8.4.3.2 char DataLink_analysis::eth_source[ETH_ALEN] [protected] Referenced by DataLink_analysis(), and get_ether_source().

8.4.3.3 unsigned short DataLink_analysis::ether_type [protected] Referenced by DataLink_analysis(), and get_ether_type(). The documentation for this class was generated from the following files:

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.h

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.cpp

8.5 DB_control Class Reference

#include <simtl_db.h>

Public Member Functions

• void close_file (int i)

• void column_plus (void)

• DB_control ()

• int get_column (void)

• FILE ∗ get_file (int i)

• QStandardItemModel ∗ get_itemModel (void)

• int get_rows (void)

• FILE ∗ get_sum (void)

• QTableView ∗ get_table (void)

• void initial (void)

• void row_plus (void)

• void set_itemModel (QObject ∗parent)

• void set_model (void)

• void set_table (QTableView ∗q)

Private Attributes

• int column

• FILE ∗∗ files

• QStandardItemModel ∗ itemModel

• QTableView ∗ qtable

• int rows

• FILE ∗ sum

8.5.1 Constructor & Destructor Documentation

8.5.1.1 DB_control::DB_control ()

8.5.2 Member Function Documentation

8.5.2.1 void DB_control::close_file (int i) [inline]

References files.

8.5.2.2 void DB_control::column_plus (void) [inline]

References column. Referenced by Arp_analysis::Arp_analysis(), Icmp_analysis::Icmp_analysis(), Ip_analysis::Ip_- analysis(), and Tcp_analysis::Tcp_analysis().

8.5 DB_control Class Reference  27

8.5.2.3 int DB_control::get_column (void) [inline] References column. Referenced by Arp_analysis::Arp_analysis(), Icmp_analysis::Icmp_analysis(), Ip_analysis::Ip_- analysis(), and Tcp_analysis::Tcp_analysis().

8.5.2.4 FILE∗ DB_control::get_file (int i) [inline] References files. Referenced by Arp_analysis::Arp_analysis(), DataLink_analysis::DataLink_analysis(), Analyzer::ether_Analysis(), Icmp_analysis::Icmp_analysis(), Ip_analysis::Ip_analysis(), and Tcp_analysis::Tcp_analysis().

8.5.2.5 QStandardItemModel∗ DB_control::get_itemModel (void) [inline] References itemModel. Referenced by Arp_analysis::Arp_analysis(), Icmp_analysis::Icmp_analysis(), Ip_analysis::Ip_- analysis(), Simtl_gui::table_clicked(), and Tcp_analysis::Tcp_analysis().

8.5.2.6 int DB_control::get_rows (void) [inline] References rows. Referenced by Arp_analysis::Arp_analysis(), Icmp_analysis::Icmp_analysis(), Ip_analysis::Ip_- analysis(), and Tcp_analysis::Tcp_analysis().

8.5.2.7 FILE∗ DB_control::get_sum (void) [inline] References sum. Referenced by Arp_analysis::Arp_analysis(), DataLink_analysis::DataLink_analysis(), Icmp_- analysis::Icmp_analysis(), Ip_analysis::Ip_analysis(), and Tcp_analysis::Tcp_analysis().

8.5.2.8 QTableView∗ DB_control::get_table (void) [inline] References qtable. Referenced by Analyzer::ether_Analysis().

8.5.2.9 void DB_control::initial (void) References column, files, Pcap_option::get_packetcnt(), p_option, rows, and sum. Referenced by Simtl_gui::button_start().

8.5.2.10 void DB_control::row_plus (void) [inline] References column, and rows. Referenced by Analyzer::ether_Analysis().

Generated on Fri May 30 23:06:20 2008 for simtl2 by Doxygen

28  Class Documentation

8.5.2.11 void DB_control::set_itemModel (QObject ∗ parent) References itemModel, and set_model(). Referenced by main().

8.5.2.12 void DB_control::set_model (void) [inline] References itemModel, and qtable. Referenced by Analyzer::ether_Analysis(), and set_itemModel().

8.5.2.13 void DB_control::set_table (QTableView ∗ q) [inline] References qtable. Referenced by main(), and Simtl_gui::Simtl_gui().

8.5.3 Member Data Documentation

8.5.3.1 int DB_control::column [private] Referenced by column_plus(), get_column(), initial(), and row_plus().

8.5.3.2 FILE∗∗ DB_control::files [private] Referenced by close_file(), get_file(), and initial().

8.5.3.3 QStandardItemModel∗ DB_control::itemModel [private] Referenced by get_itemModel(), set_itemModel(), and set_model().

8.5.3.4 QTableView∗ DB_control::qtable [private] Referenced by get_table(), set_model(), and set_table().

8.5.3.5 int DB_control::rows [private] Referenced by get_rows(), initial(), and row_plus().

8.5.3.6 FILE∗ DB_control::sum [private] Referenced by get_sum(), and initial(). The documentation for this class was generated from the following files:

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_db.h

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_db.cpp

8.6 Icmp_analysis Class Reference

#include <simtl_analysis.h>

Inheritance diagram for Icmp_analysis::

Public Member Functions

• unsigned short get_icmp_check (void)

• unsigned char get_icmp_code (void)

• unsigned char get_icmp_type (void)

• Icmp_analysis (struct ether_header ∗ep, int iph_size)

Private Attributes

• unsigned short icmp_check

• unsigned char icmp_code

• unsigned char icmp_type

• struct icmp ∗ icmph

8.6.1 Constructor & Destructor Documentation

8.6.1.1 Icmp_analysis::Icmp_analysis (struct ether_header ∗ ep, int iph_size)

References DB_control::column_plus(), db_control, Pcap_option::get_chcnt(), DB_- control::get_column(), DB_control::get_file(), DB_control::get_itemModel(), DB_- control::get_rows(), DB_control::get_sum(), icmp_check, icmp_code, icmp_type, icmph, and Packet_analysis::p_point.

8.6.2 Member Function Documentation

8.6.2.1 unsigned short Icmp_analysis::get_icmp_check (void) [inline]

References icmp_check.

8.6.2.2 unsigned char Icmp_analysis::get_icmp_code (void) [inline]

References icmp_code.

8.6.2.3 unsigned char Icmp_analysis::get_icmp_type (void) [inline]

References icmp_type.

30  Class Documentation

8.6.3 Member Data Documentation

8.6.3.1 unsigned short Icmp_analysis::icmp_check [private] Referenced by get_icmp_check(), and Icmp_analysis().

8.6.3.2 unsigned char Icmp_analysis::icmp_code [private] Referenced by get_icmp_code(), and Icmp_analysis().

8.6.3.3 unsigned char Icmp_analysis::icmp_type [private] Referenced by get_icmp_type(), and Icmp_analysis().

8.6.3.4 struct icmp∗ Icmp_analysis::icmph [read, private] Referenced by Icmp_analysis(). The documentation for this class was generated from the following files:

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.h

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.cpp

8.7 Ip_analysis Class Reference

#include <simtl_analysis.h>

Inheritance diagram for Ip_analysis::

Public Member Functions

• unsigned short get_ip_check (void)

• char ∗ get_ip_dest ()

• unsigned short get_ip_fragoff (void)

• unsigned char get_ip_hlen (void)

• unsigned short get_ip_id (void)

• unsigned char get_ip_proto (void)

• char ∗ get_ip_source ()

• unsigned char get_ip_tos (void)

• unsigned short get_ip_totlen (void)

• unsigned char get_ip_ttl (void)

• unsigned char get_ip_ver (void)

• Ip_analysis (struct iphdr ∗iphd)

• Ip_analysis (struct ether_header ∗ep)

Protected Attributes

• unsigned short ip_check

• char ip_destination [IP_ALEN]

• unsigned short ip_fragoff

• unsigned char ip_hlen

• unsigned short ip_id

• unsigned char ip_proto

• char ip_source [IP_ALEN]

• unsigned char ip_tos

• unsigned short ip_totlen

• unsigned char ip_ttl

• unsigned char ip_ver

• struct iphdr ∗ iph

• char ∗ p_point

32  Class Documentation

8.7.1 Constructor & Destructor Documentation

8.7.1.1 Ip_analysis::Ip_analysis (struct ether_header ∗ ep) References DB_control::column_plus(), db_control, Pcap_option::get_chcnt(), DB_- control::get_column(), DB_control::get_file(), DB_control::get_itemModel(), DB_- control::get_rows(), DB_control::get_sum(), Time_control::get_time(), ip_check, ip_- destination, ip_fragoff, ip_hlen, ip_id, ip_proto, ip_source, ip_tos, ip_totlen, ip_ttl, ip_ver, iph, p_point, and time_control.

8.7.1.2 Ip_analysis::Ip_analysis (struct iphdr ∗ iphd) References db_control, Pcap_option::get_chcnt(), DB_control::get_file(), DB_control::get_- sum(), ip_check, ip_destination, ip_fragoff, ip_hlen, ip_id, ip_proto, ip_source, ip_tos, ip_- totlen, ip_ttl, and ip_ver.

8.7.2 Member Function Documentation

8.7.2.1 unsigned short Ip_analysis::get_ip_check (void) [inline] References ip_check.

8.7.2.2 char∗ Ip_analysis::get_ip_dest () [inline] References ip_destination.

8.7.2.3 unsigned short Ip_analysis::get_ip_fragoff (void) [inline] References ip_fragoff.

8.7.2.4 unsigned char Ip_analysis::get_ip_hlen (void) [inline] References ip_hlen. Referenced by Analyzer::ether_Analysis().

8.7.2.5 unsigned short Ip_analysis::get_ip_id (void) [inline] References ip_id.

8.7.2.6 unsigned char Ip_analysis::get_ip_proto (void) [inline] References ip_proto. Referenced by Analyzer::ether_Analysis().

8.7.2.7 char∗ Ip_analysis::get_ip_source () [inline] References ip_source.

Generated on Fri May 30 23:06:20 2008 for simtl2 by Doxygen

8.7 Ip_analysis Class Reference  33

8.7.2.8 unsigned char Ip_analysis::get_ip_tos (void) [inline] References ip_tos.

8.7.2.9 unsigned short Ip_analysis::get_ip_totlen (void) [inline] References ip_totlen. Referenced by Analyzer::ether_Analysis().

8.7.2.10 unsigned char Ip_analysis::get_ip_ttl (void) [inline] References ip_ttl.

8.7.2.11 unsigned char Ip_analysis::get_ip_ver (void) [inline] References ip_ver.

8.7.3 Member Data Documentation

8.7.3.1 unsigned short Ip_analysis::ip_check [protected] Referenced by get_ip_check(), and Ip_analysis().

8.7.3.2 char Ip_analysis::ip_destination[IP_ALEN] [protected] Referenced by get_ip_dest(), and Ip_analysis().

8.7.3.3 unsigned short Ip_analysis::ip_fragoff [protected] Referenced by get_ip_fragoff(), and Ip_analysis().

8.7.3.4 unsigned char Ip_analysis::ip_hlen [protected] Referenced by get_ip_hlen(), and Ip_analysis().

8.7.3.5 unsigned short Ip_analysis::ip_id [protected] Referenced by get_ip_id(), and Ip_analysis().

8.7.3.6 unsigned char Ip_analysis::ip_proto [protected] Referenced by get_ip_proto(), and Ip_analysis().

8.7.3.7 char Ip_analysis::ip_source[IP_ALEN] [protected] Referenced by get_ip_source(), and Ip_analysis().

Generated on Fri May 30 23:06:20 2008 for simtl2 by Doxygen

34  Class Documentation

8.7.3.8 unsigned char Ip_analysis::ip_tos [protected] Referenced by get_ip_tos(), and Ip_analysis().

8.7.3.9 unsigned short Ip_analysis::ip_totlen [protected] Referenced by get_ip_totlen(), and Ip_analysis().

8.7.3.10 unsigned char Ip_analysis::ip_ttl [protected] Referenced by get_ip_ttl(), and Ip_analysis().

8.7.3.11 unsigned char Ip_analysis::ip_ver [protected] Referenced by get_ip_ver(), and Ip_analysis().

8.7.3.12 struct iphdr∗ Ip_analysis::iph [read, protected] Referenced by Ip_analysis().

8.7.3.13 char∗ Ip_analysis::p_point [protected] Reimplemented from Packet_analysis (p. 35). Referenced by Ip_analysis(). The documentation for this class was generated from the following files:

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.h

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.cpp

8.8 Packet_analysis Class Reference  35

8.8 Packet_analysis Class Reference

#include <simtl_analysis.h>

Inheritance diagram for Packet_analysis::

Public Member Functions

• void data_analysis (struct ether_header ∗ep, int totlen, FILE ∗file)

Public Attributes

• int i

• char ∗ p_point

8.8.1 Member Function Documentation

8.8.1.1 void Packet_analysis::data_analysis (struct ether_header ∗ ep, int totlen,

FILE ∗ file)

Referenced by Analyzer::ether_Analysis().

8.8.2 Member Data Documentation

8.8.2.1 int Packet_analysis::i Referenced by DataLink_analysis::DataLink_analysis().

8.8.2.2 char∗ Packet_analysis::p_point Reimplemented in Ip_analysis (p. 34), and Tcp_analysis (p. 45). Referenced by Arp_analysis::Arp_analysis(), and Icmp_analysis::Icmp_analysis(). The documentation for this class was generated from the following files:

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.h

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.cpp

8.9 Pcap_option Class Reference

#include <simtl_analysis.h>

Public Member Functions

• bool get_cflag (void)

• int get_chcnt (void)

• char ∗ get_device (void)

• int get_packetcnt (void)

• pcap_t ∗ get_pcap_t (void)

• char ∗ get_rule (void)

• Pcap_option ()

• void plus_chcnt (void)

• void set_cflag (bool flag)

• void set_chcnt (int cnt)

• void set_device (char ∗dev)

• void set_packetcnt (int cnt)

• void set_pcap_t (pcap_t ∗p)

• void set_rule (char ∗rule)

Private Attributes

• bool cflag

• int chcnt

• char device [10]

• char filter_rule [100]

• int packetcnt

• pcap_t ∗ pd

8.9.1 Constructor & Destructor Documentation

8.9.1.1 Pcap_option::Pcap_option () [inline]

References cflag, chcnt, device, filter_rule, and packetcnt.

8.9.2 Member Function Documentation

8.9.2.1 bool Pcap_option::get_cflag (void) [inline]

References cflag.

8.9.2.2 int Pcap_option::get_chcnt (void) [inline]

References chcnt.

Referenced by Arp_analysis::Arp_analysis(), DataLink_analysis::DataLink_analysis(), Analyzer::ether_Analysis(), Icmp_analysis::Icmp_analysis(), Ip_analysis::Ip_analysis(), and Tcp_analysis::Tcp_analysis().

8.9 Pcap_option Class Reference  37

8.9.2.3 char∗ Pcap_option::get_device (void) [inline] References device. Referenced by init_Analysis().

8.9.2.4 int Pcap_option::get_packetcnt (void) [inline] References packetcnt. Referenced by init_Analysis(), and DB_control::initial().

8.9.2.5 pcap_t∗ Pcap_option::get_pcap_t (void) [inline] References pd. Referenced by init_Analysis().

8.9.2.6 char∗ Pcap_option::get_rule (void) [inline] References filter_rule. Referenced by init_Analysis(), and Simtl_gui::set_options().

8.9.2.7 void Pcap_option::plus_chcnt (void) [inline] References chcnt. Referenced by Analyzer::ether_Analysis().

8.9.2.8 void Pcap_option::set_cflag (bool flag) [inline] References cflag.

8.9.2.9 void Pcap_option::set_chcnt (int cnt) [inline] References chcnt. Referenced by Simtl_gui::button_start().

8.9.2.10 void Pcap_option::set_device (char ∗ dev) [inline] References device. Referenced by init_Analysis().

8.9.2.11 void Pcap_option::set_packetcnt (int cnt) [inline] References packetcnt. Referenced by Simtl_gui::set_options().

Generated on Fri May 30 23:06:20 2008 for simtl2 by Doxygen

38  Class Documentation

8.9.2.12 void Pcap_option::set_pcap_t (pcap_t ∗ p) [inline] References pd. Referenced by init_Analysis().

8.9.2.13 void Pcap_option::set_rule (char ∗ rule) [inline] References filter_rule. Referenced by Simtl_gui::set_options().

8.9.3 Member Data Documentation

8.9.3.1 bool Pcap_option::cflag [private] Referenced by get_cflag(), Pcap_option(), and set_cflag().

8.9.3.2 int Pcap_option::chcnt [private] Referenced by get_chcnt(), Pcap_option(), plus_chcnt(), and set_chcnt().

8.9.3.3 char Pcap_option::device[10] [private] Referenced by get_device(), Pcap_option(), and set_device().

8.9.3.4 char Pcap_option::filter_rule[100] [private] Referenced by get_rule(), Pcap_option(), and set_rule().

8.9.3.5 int Pcap_option::packetcnt [private] Referenced by get_packetcnt(), Pcap_option(), and set_packetcnt().

8.9.3.6 pcap_t∗ Pcap_option::pd [private] Referenced by get_pcap_t(), and set_pcap_t(). The documentation for this class was generated from the following file:

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.h

8.10 Simtl_dir Class Reference

#include <simtl_dir.h>

Public Slots

• int changeDir ()

• int itemSelected ()

• int makeDir ()

• int okDir ()

• int removeDir ()

• int renameDir ()

Public Member Functions

• Simtl_dir (QWidget ∗parent=0)

Private Attributes

• QDir ∗ directory

• QListWidget ∗ dirlist

• QLineEdit ∗ filename

8.10.1 Constructor & Destructor Documentation

8.10.1.1 Simtl_dir::Simtl_dir (QWidget ∗ parent = 0)

References changeDir(), directory, dirlist, filename, itemSelected(), makeDir(), okDir(), removeDir(), and renameDir().

8.10.2 Member Function Documentation

8.10.2.1 int Simtl_dir::changeDir () [slot]

References directory, and dirlist.

Referenced by Simtl_dir().

8.10.2.2 int Simtl_dir::itemSelected () [slot]

References dirlist, and filename.

Referenced by Simtl_dir().

8.10.2.3 int Simtl_dir::makeDir () [slot]

References directory, dirlist, and filename.

Referenced by Simtl_dir().

40  Class Documentation

8.10.2.4 int Simtl_dir::okDir () [slot] References directory, and dirlist. Referenced by Simtl_dir().

8.10.2.5 int Simtl_dir::removeDir () [slot] References directory, dirlist, and filename. Referenced by Simtl_dir().

8.10.2.6 int Simtl_dir::renameDir () [slot] References directory, dirlist, and filename. Referenced by Simtl_dir().

8.10.3 Member Data Documentation

8.10.3.1 QDir∗ Simtl_dir::directory [private] Referenced by changeDir(), makeDir(), okDir(), removeDir(), renameDir(), and Simtl_dir().

8.10.3.2 QListWidget∗ Simtl_dir::dirlist [private] Referenced by changeDir(), itemSelected(), makeDir(), okDir(), removeDir(), renameDir(), and Simtl_dir().

8.10.3.3 QLineEdit∗ Simtl_dir::filename [private] Referenced by itemSelected(), makeDir(), removeDir(), renameDir(), and Simtl_dir(). The documentation for this class was generated from the following files:

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_dir.h

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_dir.cpp

8.11 Simtl_gui Class Reference  41

8.11 Simtl_gui Class Reference

#include <simtl_gui.h>

Public Member Functions

• QTableView ∗ get_table (void)

• Simtl_gui (QMainWindow ∗parent=0)

• void table_resize (void)

Private Slots

• void button_save (void)

• void button_start (void)

• void set_options (int number)

• void set_options (QString str)

• void table_clicked (QModelIndex lindex)

Private Attributes

• Simtl_dir ∗ simtl_dir

• Ui::MainWindow ui

8.11.1 Constructor & Destructor Documentation

8.11.1.1 Simtl_gui::Simtl_gui (QMainWindow ∗ parent = 0) References button_save(), button_start(), db_control, Ui_MainWindow::filterEdit, Ui_- MainWindow::packetBox, Ui_MainWindow::saveButton, set_options(), DB_control::set_- table(), Ui_MainWindow::setupUi(), Ui_MainWindow::startButton, Ui_MainWindow::sumView, table_clicked(), and ui.

8.11.2 Member Function Documentation

8.11.2.1 void Simtl_gui::button_save (void) [private, slot] References simtl_dir. Referenced by Simtl_gui().

8.11.2.2 void Simtl_gui::button_start (void) [private, slot] References db_control, init_Analysis(), DB_control::initial(), p_option, Pcap_option::set_- chcnt(), Time_control::start_clock(), and time_control. Referenced by Simtl_gui().

Generated on Fri May 30 23:06:20 2008 for simtl2 by Doxygen

42  Class Documentation

8.11.2.3 QTableView∗ Simtl_gui::get_table (void) [inline] References Ui_MainWindow::sumView, and ui. Referenced by main().

8.11.2.4 void Simtl_gui::set_options (int number) [private, slot] References p_option, and Pcap_option::set_packetcnt().

8.11.2.5 void Simtl_gui::set_options (QString str) [private, slot] References Pcap_option::get_rule(), p_option, and Pcap_option::set_rule(). Referenced by Simtl_gui().

8.11.2.6 void Simtl_gui::table_clicked (QModelIndex lindex) [private, slot] References db_control, Ui_MainWindow::detailBrowser, DB_control::get_itemModel(), and ui. Referenced by Simtl_gui().

8.11.2.7 void Simtl_gui::table_resize (void) References Ui_MainWindow::sumView, and ui. Referenced by main().

8.11.3 Member Data Documentation

8.11.3.1 Simtl_dir∗ Simtl_gui::simtl_dir [private] Referenced by button_save().

8.11.3.2 Ui::MainWindow Simtl_gui::ui [private] Referenced by get_table(), Simtl_gui(), table_clicked(), and table_resize(). The documentation for this class was generated from the following files:

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_gui.h

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_gui.cpp

8.12 Tcp_analysis Class Reference

#include <simtl_analysis.h>

Inheritance diagram for Tcp_analysis::

Public Member Functions

• bool get_tcp_ack (void)

• unsigned int get_tcp_anumber (void)

• unsigned short get_tcp_check (void)

• unsigned char get_tcp_doff (void)

• unsigned short get_tcp_dport (void)

• bool get_tcp_fin (void)

• bool get_tcp_psh (void)

• bool get_tcp_rst (void)

• unsigned int get_tcp_snumber (void)

• unsigned short get_tcp_sport (void)

• bool get_tcp_syn (void)

• unsigned short get_tcp_upoint (void)

• bool get_tcp_urg (void)

• unsigned short get_tcp_wins (void)

• Tcp_analysis (struct ether_header ∗ep, int iph_size)

Protected Attributes

• char ∗ p_point

• bool tcp_ack

• unsigned int tcp_anumber

• unsigned short tcp_check

• unsigned char tcp_doff

• unsigned short tcp_dport

• bool tcp_fin

• bool tcp_psh

• bool tcp_rst

• unsigned int tcp_snumber

• unsigned short tcp_sport

• bool tcp_syn

• unsigned short tcp_upoint

• bool tcp_urg

• unsigned short tcp_wins

• struct tcphdr ∗ tcph

44  Class Documentation

8.12.1 Constructor & Destructor Documentation

8.12.1.1 Tcp_analysis::Tcp_analysis (struct ether_header ∗ ep, int iph_size) References DB_control::column_plus(), db_control, Pcap_option::get_chcnt(), DB_- control::get_column(), DB_control::get_file(), DB_control::get_itemModel(), DB_- control::get_rows(), DB_control::get_sum(), p_point, tcp_ack, tcp_anumber, tcp_check, tcp_doff, tcp_dport, tcp_fin, tcp_psh, tcp_rst, tcp_snumber, tcp_sport, tcp_syn, tcp_upoint, tcp_urg, tcp_wins, and tcph.

8.12.2 Member Function Documentation

8.12.2.1 bool Tcp_analysis::get_tcp_ack (void) [inline] References tcp_ack.

8.12.2.2 unsigned int Tcp_analysis::get_tcp_anumber (void) [inline] References tcp_anumber.

8.12.2.3 unsigned short Tcp_analysis::get_tcp_check (void) [inline] References tcp_check.

8.12.2.4 unsigned char Tcp_analysis::get_tcp_doff (void) [inline] References tcp_doff.

8.12.2.5 unsigned short Tcp_analysis::get_tcp_dport (void) [inline] References tcp_dport.

8.12.2.6 bool Tcp_analysis::get_tcp_fin (void) [inline] References tcp_fin.

8.12.2.7 bool Tcp_analysis::get_tcp_psh (void) [inline] References tcp_psh.

8.12.2.8 bool Tcp_analysis::get_tcp_rst (void) [inline] References tcp_rst.

8.12.2.9 unsigned int Tcp_analysis::get_tcp_snumber (void) [inline] References tcp_snumber.

Generated on Fri May 30 23:06:20 2008 for simtl2 by Doxygen

8.12 Tcp_analysis Class Reference  45

8.12.2.10 unsigned short Tcp_analysis::get_tcp_sport (void) [inline] References tcp_sport.

8.12.2.11 bool Tcp_analysis::get_tcp_syn (void) [inline] References tcp_syn.

8.12.2.12 unsigned short Tcp_analysis::get_tcp_upoint (void) [inline] References tcp_upoint.

8.12.2.13 bool Tcp_analysis::get_tcp_urg (void) [inline] References tcp_urg.

8.12.2.14 unsigned short Tcp_analysis::get_tcp_wins (void) [inline] References tcp_wins.

8.12.3 Member Data Documentation

8.12.3.1 char∗ Tcp_analysis::p_point [protected] Reimplemented from Packet_analysis (p. 35). Referenced by Tcp_analysis().

8.12.3.2 bool Tcp_analysis::tcp_ack [protected] Referenced by get_tcp_ack(), and Tcp_analysis().

8.12.3.3 unsigned int Tcp_analysis::tcp_anumber [protected] Referenced by get_tcp_anumber(), and Tcp_analysis().

8.12.3.4 unsigned short Tcp_analysis::tcp_check [protected] Referenced by get_tcp_check(), and Tcp_analysis().

8.12.3.5 unsigned char Tcp_analysis::tcp_doff [protected] Referenced by get_tcp_doff(), and Tcp_analysis().

8.12.3.6 unsigned short Tcp_analysis::tcp_dport [protected] Referenced by get_tcp_dport(), and Tcp_analysis().

Generated on Fri May 30 23:06:20 2008 for simtl2 by Doxygen

46  Class Documentation

8.12.3.7 bool Tcp_analysis::tcp_fin [protected] Referenced by get_tcp_fin(), and Tcp_analysis().

8.12.3.8 bool Tcp_analysis::tcp_psh [protected] Referenced by get_tcp_psh(), and Tcp_analysis().

8.12.3.9 bool Tcp_analysis::tcp_rst [protected] Referenced by get_tcp_rst(), and Tcp_analysis().

8.12.3.10 unsigned int Tcp_analysis::tcp_snumber [protected] Referenced by get_tcp_snumber(), and Tcp_analysis().

8.12.3.11 unsigned short Tcp_analysis::tcp_sport [protected] Referenced by get_tcp_sport(), and Tcp_analysis().

8.12.3.12 bool Tcp_analysis::tcp_syn [protected] Referenced by get_tcp_syn(), and Tcp_analysis().

8.12.3.13 unsigned short Tcp_analysis::tcp_upoint [protected] Referenced by get_tcp_upoint(), and Tcp_analysis().

8.12.3.14 bool Tcp_analysis::tcp_urg [protected] Referenced by get_tcp_urg(), and Tcp_analysis().

8.12.3.15 unsigned short Tcp_analysis::tcp_wins [protected] Referenced by get_tcp_wins(), and Tcp_analysis().

8.12.3.16 struct tcphdr∗ Tcp_analysis::tcph [read, protected] Referenced by Tcp_analysis(). The documentation for this class was generated from the following files:

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.h

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.cpp

8.13 Time_control Class Reference  47

8.13 Time_control Class Reference

#include <simtl_time.h>

Public Member Functions

• char ∗ get_time (void)

• void start_clock (void)

• Time_control (void)

Private Attributes

• struct timeval cur_time

• struct timeval start_time

8.13.1 Constructor & Destructor Documentation

8.13.1.1 Time_control::Time_control (void)

8.13.2 Member Function Documentation

8.13.2.1 char ∗ Time_control::get_time (void) References cur_time, and start_time. Referenced by Arp_analysis::Arp_analysis(), and Ip_analysis::Ip_analysis().

8.13.2.2 void Time_control::start_clock (void) References start_time. Referenced by Simtl_gui::button_start().

8.13.3 Member Data Documentation

8.13.3.1 struct timeval Time_control::cur_time [read, private] Referenced by get_time().

8.13.3.2 struct timeval Time_control::start_time [read, private] Referenced by get_time(), and start_clock(). The documentation for this class was generated from the following files:

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_time.h

• Desktop/Study/Programming/KDEProject/simtl2/src/simtl_time.cpp

48  Class Documentation

8.14 Ui::MainWindow Class Reference

#include <ui_simtl_gui.h>

Inheritance diagram for Ui::MainWindow::

The documentation for this class was generated from the following file:

• Desktop/Study/Programming/KDEProject/simtl2/src/ui_simtl_gui.h

8.15 Ui_MainWindow Class Reference

#include <ui_simtl_gui.h>

Inheritance diagram for Ui_MainWindow::

Public Member Functions

• void retranslateUi (QMainWindow ∗MainWindow)

• void setupUi (QMainWindow ∗MainWindow)

Public Attributes

• QPushButton ∗ cancelButton

• QWidget ∗ centralwidget

• QTextEdit ∗ detailBrowser

• QLineEdit ∗ filterEdit

• QHBoxLayout ∗ hboxLayout

• QHBoxLayout ∗ hboxLayout1

• QHBoxLayout ∗ hboxLayout2

• QHBoxLayout ∗ hboxLayout3

• QLabel ∗ label

• QLabel ∗ label_2

• QWidget ∗ layoutWidget

• QWidget ∗ layoutWidget1

• QWidget ∗ layoutWidget2

• QMenuBar ∗ menubar

• QSpinBox ∗ packetBox

• QPushButton ∗ saveButton

• QPushButton ∗ startButton

• QStatusBar ∗ statusbar

• QPushButton ∗ stopButton

• QTableView ∗ sumView

• QToolBar ∗ toolBar

• QWidget ∗ widget

8.15.1 Member Function Documentation

8.15.1.1 void Ui_MainWindow::retranslateUi (QMainWindow ∗ MainWindow)

       [inline]

References cancelButton, label, label_2, saveButton, startButton, stopButton, and toolBar.

Referenced by setupUi().

50  Class Documentation

8.15.1.2 void Ui_MainWindow::setupUi (QMainWindow ∗ MainWindow)

          [inline]

References cancelButton, centralwidget, detailBrowser, filterEdit, hboxLayout, hboxLayout1, hboxLayout2, hboxLayout3, label, label_2, layoutWidget, layoutWidget1, layoutWidget2, menubar, packetBox, retranslateUi(), saveButton, startButton, statusbar, stopButton, sumView, toolBar, and widget. Referenced by Simtl_gui::Simtl_gui().

8.15.2 Member Data Documentation

8.15.2.1 QPushButton∗ Ui_MainWindow::cancelButton Referenced by retranslateUi(), and setupUi().

8.15.2.2 QWidget∗ Ui_MainWindow::centralwidget Referenced by setupUi().

8.15.2.3 QTextEdit∗ Ui_MainWindow::detailBrowser Referenced by setupUi(), and Simtl_gui::table_clicked().

8.15.2.4 QLineEdit∗ Ui_MainWindow::filterEdit Referenced by setupUi(), and Simtl_gui::Simtl_gui().

8.15.2.5 QHBoxLayout∗ Ui_MainWindow::hboxLayout Referenced by setupUi().

8.15.2.6 QHBoxLayout∗ Ui_MainWindow::hboxLayout1 Referenced by setupUi().

8.15.2.7 QHBoxLayout∗ Ui_MainWindow::hboxLayout2 Referenced by setupUi().

8.15.2.8 QHBoxLayout∗ Ui_MainWindow::hboxLayout3 Referenced by setupUi().

8.15.2.9 QLabel∗ Ui_MainWindow::label Referenced by retranslateUi(), and setupUi().

Generated on Fri May 30 23:06:20 2008 for simtl2 by Doxygen

8.15 Ui_MainWindow Class Reference  51

8.15.2.10 QLabel∗ Ui_MainWindow::label_2 Referenced by retranslateUi(), and setupUi().

8.15.2.11 QWidget∗ Ui_MainWindow::layoutWidget Referenced by setupUi().

8.15.2.12 QWidget∗ Ui_MainWindow::layoutWidget1 Referenced by setupUi().

8.15.2.13 QWidget∗ Ui_MainWindow::layoutWidget2 Referenced by setupUi().

8.15.2.14 QMenuBar∗ Ui_MainWindow::menubar Referenced by setupUi().

8.15.2.15 QSpinBox∗ Ui_MainWindow::packetBox Referenced by setupUi(), and Simtl_gui::Simtl_gui().

8.15.2.16 QPushButton∗ Ui_MainWindow::saveButton Referenced by retranslateUi(), setupUi(), and Simtl_gui::Simtl_gui().

8.15.2.17 QPushButton∗ Ui_MainWindow::startButton Referenced by retranslateUi(), setupUi(), and Simtl_gui::Simtl_gui().

8.15.2.18 QStatusBar∗ Ui_MainWindow::statusbar Referenced by setupUi().

8.15.2.19 QPushButton∗ Ui_MainWindow::stopButton Referenced by retranslateUi(), and setupUi().

8.15.2.20 QTableView∗ Ui_MainWindow::sumView Referenced by Simtl_gui::get_table(), setupUi(), Simtl_gui::Simtl_gui(), and Simtl_gui::table_- resize().

Generated on Fri May 30 23:06:20 2008 for simtl2 by Doxygen

52  Class Documentation

8.15.2.21 QToolBar∗ Ui_MainWindow::toolBar Referenced by retranslateUi(), and setupUi().

8.15.2.22 QWidget∗ Ui_MainWindow::widget Referenced by setupUi(). The documentation for this class was generated from the following file:

• Desktop/Study/Programming/KDEProject/simtl2/src/ui_simtl_gui.h

Chapter 9

File Documentation

9.1 Desktop/Study/Programming/KDEProject/simtl2/src/analyzer_-

interface.h File Reference

Classes

• class AnalyzerInterface

9.2 Desktop/Study/Programming/KDEProject/simtl2/src/main.cpp File Reference

#include <QApplication>
#include "simtl_analysis.h"
#include "simtl_db.h"
#include "simtl_gui.h"
#include "ui_simtl_gui.h"

Functions

• int main (int argc, char ∗argv[ ])

9.2.1 Function Documentation

9.2.1.1 int main (int argc, char ∗ argv[ ])

References db_control, Simtl_gui::get_table(), DB_control::set_itemModel(), DB_- control::set_table(), and Simtl_gui::table_resize().

9.3 Desktop/Study/Programming/KDEProject/simtl2/src/moc_simtl_dir.cpp File

9.3 Desktop/Study/Programming/KDEProject/simtl2/src/moc_- simtl_dir.cpp File Reference

#include "simtl_dir.h"

Variables

• static const uint qt_meta_data_Simtl_dir [ ]

• static const char qt_meta_stringdata_Simtl_dir [ ]

9.3.1 Variable Documentation

9.3.1.1 const uint qt_meta_data_Simtl_dir[ ] [static]

Initial value:

 {

       1,
       0,
       0,    0,
       6,   10,
       0,    0,
       0,    0,

      15,   10,   11,   10, 0x0a,
      30,   10,   11,   10, 0x0a,
      42,   10,   11,   10, 0x0a,
      52,   10,   11,   10, 0x0a,
      64,   10,   11,   10, 0x0a,
      76,   10,   11,   10, 0x0a,
       0
}

9.3.1.2 const char qt_meta_stringdata_Simtl_dir[ ] [static]

Initial value:

 {
    "Simtl_dir\0\0int\0itemSelected()\0changeDir()\0"
    "makeDir()\0removeDir()\0renameDir()\0"
    "okDir()\0"
}

9.4 Desktop/Study/Programming/KDEProject/simtl2/src/moc_- simtl_gui.cpp File Reference

#include "simtl_gui.h"

Variables

• static const uint qt_meta_data_Simtl_gui [ ]

• static const char qt_meta_stringdata_Simtl_gui [ ]

9.4.1 Variable Documentation

9.4.1.1 const uint qt_meta_data_Simtl_gui[ ] [static]

Initial value:

 {

       1,
       0,
       0,    0,
       5,   10,
       0,    0,
       0,    0,

      15,   11,   10,   10, 0x08,
      43,   36,   10,   10, 0x08,
      60,   10,   10,   10, 0x08,
      75,   10,   10,   10, 0x08,
      96,   89,   10,   10, 0x08,
       0
}

9.4.1.2 const char qt_meta_stringdata_Simtl_gui[ ] [static]

Initial value:

 {
    "Simtl_gui\0\0str\0set_options(QString)\0"
    "number\0set_options(int)\0button_start()\0"
    "button_save()\0lindex\0table_clicked(QModelIndex)\0"
}

9.5 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.cpp File Reference 57

9.5 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_-

analysis.cpp File Reference

#include "simtl_analysis.h"

Defines

• #define uint32_t u_int32_t

Functions

• void init_Analysis (void)

Variables

• static Pcap_option p_option

9.5.1 Define Documentation

9.5.1.1 #define uint32_t u_int32_t

9.5.2 Function Documentation

9.5.2.1 void init_Analysis (void) References Analyzer::ether_Analysis(), Pcap_option::get_device(), Pcap_option::get_- packetcnt(), Pcap_option::get_pcap_t(), Pcap_option::get_rule(), Pcap_option::set_device(), and Pcap_option::set_pcap_t(). Referenced by Simtl_gui::button_start().

9.5.3 Variable Documentation

9.5.3.1 Pcap_option p_option [static] Referenced by Simtl_gui::button_start(), DB_control::initial(), and Simtl_gui::set_options().

9.6 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_- analysis.h File Reference

#include <sys/types.h>
#include <sys/socket.h>
#include <unistd.h>
#include <string.h>
#include <stdlib.h>
#include <pcap.h>
#include <arpa/inet.h>
#include <net/ethernet.h>
#include <net/if_arp.h>
#include <netinet/in.h>
#include <netinet/ip.h>
#include <netinet/tcp.h>
#include <netinet/ip_icmp.h>
#include "simtl_db.h"
#include "simtl_gui.h"
#include "simtl_time.h"

Classes

• class Analyzer

• class Arp_analysis

• class DataLink_analysis

• class Icmp_analysis

• class Ip_analysis

• class Packet_analysis

• class Pcap_option

• class Tcp_analysis

Defines

• #define IP_ALEN 16

Functions

• void init_Analysis (void)

Variables

• Pcap_option p_option 9.6 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_analysis.h File Reference

59

9.6.1 Define Documentation

9.6.1.1 #define IP_ALEN 16

9.6.2 Function Documentation

9.6.2.1 void init_Analysis (void) References Analyzer::ether_Analysis(), Pcap_option::get_device(), Pcap_option::get_- packetcnt(), Pcap_option::get_pcap_t(), Pcap_option::get_rule(), Pcap_option::set_device(), and Pcap_option::set_pcap_t(). Referenced by Simtl_gui::button_start().

9.6.3 Variable Documentation

9.6.3.1 Pcap_option p_option

9.7 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_- db.cpp File Reference

#include "simtl_db.h"
#include "simtl_time.h"

Variables

• static DB_control db_control

9.7.1 Variable Documentation

9.7.1.1 DB_control db_control [static]

Referenced by Arp_analysis::Arp_analysis(), Simtl_gui::button_start(), DataLink_- analysis::DataLink_analysis(), Analyzer::ether_Analysis(), Icmp_analysis::Icmp_analysis(), Ip_analysis::Ip_analysis(), main(), Simtl_gui::Simtl_gui(), Simtl_gui::table_clicked(), and Tcp_analysis::Tcp_analysis().

9.8 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_db.h File Reference

61

9.8 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_- db.h File Reference

#include <QStandardItemModel>
#include <QTableView>
#include <stdio.h>
#include <sys/stat.h>
#include <sys/types.h>
#include <fcntl.h>
#include <unistd.h>
#include "simtl_analysis.h"

Classes

• class DB_control

Variables

• DB_control db_control

9.8.1 Variable Documentation

9.8.1.1 DB_control db_control

9.9 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_- dir.cpp File Reference

#include <QFileInfo>
#include <QPushButton>
#include <QVBoxLayout>
#include <stdlib.h>
#include "simtl_dir.h"

9.10 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_dir.h File Reference

63

9.10 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_- dir.h File Reference

#include <QWidget>
#include <QDir>
#include <QListWidget>
#include <QLineEdit>

Classes

• class Simtl_dir

9.11 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_- gui.cpp File Reference

#include "simtl_gui.h"

9.12 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_gui.h File Reference

65

9.12 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_- gui.h File Reference

#include <QMainWindow>
#include <QStandardItemModel>
#include <QStandardItem>
#include <QFile>
#include "ui_simtl_gui.h"
#include "simtl_db.h"
#include "simtl_analysis.h"
#include "simtl_dir.h"

Classes

• class Simtl_gui

66  File Documentation

9.13 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_-

time.cpp File Reference

#include "simtl_time.h"

Variables

• static Time_control time_control

9.13.1 Variable Documentation

9.13.1.1 Time_control time_control [static] Referenced by Arp_analysis::Arp_analysis(), Simtl_gui::button_start(), and Ip_analysis::Ip_- analysis().

9.14 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_time.h File Reference

67

9.14 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_-

time.h File Reference

#include <stdio.h>
#include <sys/time.h>

Classes

• class Time_control

Variables

• Time_control time_control

9.14.1 Variable Documentation

9.14.1.1 Time_control time_control

9.15 Desktop/Study/Programming/KDEProject/simtl2/src/ui_- simtl_gui.h File Reference

#include <QtCore/QVariant>
#include <QtGui/QAction>
#include <QtGui/QApplication>
#include <QtGui/QButtonGroup>
#include <QtGui/QHBoxLayout>
#include <QtGui/QLabel>
#include <QtGui/QLineEdit>
#include <QtGui/QMainWindow>
#include <QtGui/QMenuBar>
#include <QtGui/QPushButton>
#include <QtGui/QSpinBox>
#include <QtGui/QStatusBar>
#include <QtGui/QTableView>
#include <QtGui/QTextEdit>
#include <QtGui/QToolBar>
#include <QtGui/QWidget>

Namespaces

• namespace Ui

Classes

• class Ui::MainWindow

• class Ui_MainWindow

Index

∼AnalyzerInterface AnalyzerInterface, 20

Analyzer, 19

ether_Analysis, 19 AnalyzerInterface, 20

∼AnalyzerInterface, 20 AnalyzerInterface, 20 operator=, 20

Arp_analysis, 21 Arp_analysis, 21 arp_hwl, 22 arp_hwt, 22 arp_op, 22 arp_prl, 23 arp_prt, 23 arp_send_haddr, 23 arp_send_paddr, 23 arp_target_haddr, 23 arp_target_paddr, 23 Arp_analysis, 21 arph, 23 get_arp_hwl, 22 get_arp_hwt, 22 get_arp_op, 22 get_arp_prl, 22 get_arp_prt, 22 get_arp_send_haddr, 22 get_arp_send_paddr, 22 get_arp_target_haddr, 22 get_arp_target_paddr, 22

arp_hwl

Arp_analysis, 22 arp_hwt

Arp_analysis, 22 arp_op

Arp_analysis, 22 arp_prl

Arp_analysis, 23 arp_prt

Arp_analysis, 23 arp_send_haddr

Arp_analysis, 23 arp_send_paddr

Arp_analysis, 23 arp_target_haddr

Arp_analysis, 23 arp_target_paddr

Arp_analysis, 23 arph

Arp_analysis, 23

button_save

Simtl_gui, 41 button_start

Simtl_gui, 41

cancelButton

Ui_MainWindow, 50 centralwidget

Ui_MainWindow, 50 cflag

Pcap_option, 38 changeDir

Simtl_dir, 39 chcnt

Pcap_option, 38 close_file

DB_control, 26 column

DB_control, 28 column_plus

DB_control, 26 cur_time

Time_control, 47

data_analysis

Packet_analysis, 35 DataLink_analysis, 24

DataLink_analysis, 24 DataLink_analysis, 24 eth_destination, 25 eth_source, 25 ether_type, 25 get_ether_dest, 24 get_ether_source, 24 get_ether_type, 24

\underline{70 INDEX}

DB_control, 26 close_file, 26 column, 28 column_plus, 26 DB_control, 26 DB_control, 26 files, 28 get_column, 26 get_file, 27 get_itemModel, 27 get_rows, 27 get_sum, 27 get_table, 27 initial, 27 itemModel, 28 qtable, 28 row_plus, 27 rows, 28 set_itemModel, 27 set_model, 28 set_table, 28 sum, 28

db_control simtl_db.cpp, 60 simtl_db.h, 61

Desktop/ Directory Reference, 11 Desktop/Study/ Directory Reference, 16 Desktop/Study/Programming/ Directory Ref-

erence, 13 Desktop/Study/Programming/KDEProject/

Directory Reference, 12 Desktop/Study/Programming/KDEProject/simtl2/

Directory Reference, 14 Desktop/Study/Programming/KDEProject/simtl2/src/

Directory Reference, 15 Desktop/Study/Programming/KDEProject/simtl2/src/analyzer_-

interface.h, 53 Desktop/Study/Programming/KDEProject/simtl2/src/main.cpp,

54 Desktop/Study/Programming/KDEProject/simtl2/src/moc_-

simtl_dir.cpp, 55 Desktop/Study/Programming/KDEProject/simtl2/src/moc_-

simtl_gui.cpp, 56 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_-

analysis.cpp, 57 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_-

analysis.h, 58 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_-

db.cpp, 60 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_-

db.h, 61 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_-

dir.cpp, 62 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_-

dir.h, 63 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_-

gui.cpp, 64 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_-

gui.h, 65 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_-

time.cpp, 66 Desktop/Study/Programming/KDEProject/simtl2/src/simtl_-

time.h, 67 Desktop/Study/Programming/KDEProject/simtl2/src/ui_-

simtl_gui.h, 68 detailBrowser

Ui_MainWindow, 50 device

Pcap_option, 38 directory

Simtl_dir, 40 dirlist

Simtl_dir, 40

eth_destination

DataLink_analysis, 25 eth_source

DataLink_analysis, 25 ether_Analysis

Analyzer, 19 ether_type

DataLink_analysis, 25

filename

Simtl_dir, 40 files

DB_control, 28 filter_rule

Pcap_option, 38 filterEdit

Ui_MainWindow, 50

get_arp_hwl

Arp_analysis, 22 get_arp_hwt

Arp_analysis, 22 get_arp_op

Arp_analysis, 22 get_arp_prl

Arp_analysis, 22 get_arp_prt

Arp_analysis, 22 get_arp_send_haddr

Arp_analysis, 22 get_arp_send_paddr

Arp_analysis, 22 get_arp_target_haddr

Arp_analysis, 22 get_arp_target_paddr

Arp_analysis, 22 get_cflag

\underline{INDEX 71}

Pcap_option, 36 get_chcnt

Pcap_option, 36 get_column

DB_control, 26 get_device

Pcap_option, 36 get_ether_dest

DataLink_analysis, 24 get_ether_source

DataLink_analysis, 24 get_ether_type

DataLink_analysis, 24 get_file

DB_control, 27 get_icmp_check

Icmp_analysis, 29 get_icmp_code

Icmp_analysis, 29 get_icmp_type

Icmp_analysis, 29 get_ip_check

Ip_analysis, 32 get_ip_dest

Ip_analysis, 32 get_ip_fragoff

Ip_analysis, 32 get_ip_hlen

Ip_analysis, 32 get_ip_id

Ip_analysis, 32 get_ip_proto

Ip_analysis, 32 get_ip_source

Ip_analysis, 32 get_ip_tos

Ip_analysis, 32 get_ip_totlen

Ip_analysis, 33 get_ip_ttl

Ip_analysis, 33 get_ip_ver

Ip_analysis, 33 get_itemModel

DB_control, 27 get_packetcnt

Pcap_option, 37 get_pcap_t

Pcap_option, 37 get_rows

DB_control, 27 get_rule

Pcap_option, 37 get_sum

DB_control, 27 get_table

DB_control, 27 Simtl_gui, 41

get_tcp_ack

Tcp_analysis, 44 get_tcp_anumber

Tcp_analysis, 44 get_tcp_check

Tcp_analysis, 44 get_tcp_doff

Tcp_analysis, 44 get_tcp_dport

Tcp_analysis, 44 get_tcp_fin

Tcp_analysis, 44 get_tcp_psh

Tcp_analysis, 44 get_tcp_rst

Tcp_analysis, 44 get_tcp_snumber

Tcp_analysis, 44 get_tcp_sport

Tcp_analysis, 44 get_tcp_syn

Tcp_analysis, 45 get_tcp_upoint

Tcp_analysis, 45 get_tcp_urg

Tcp_analysis, 45 get_tcp_wins

Tcp_analysis, 45 get_time

Time_control, 47

hboxLayout

Ui_MainWindow, 50 hboxLayout1

Ui_MainWindow, 50 hboxLayout2

Ui_MainWindow, 50 hboxLayout3

Ui_MainWindow, 50

i

Packet_analysis, 35 Icmp_analysis, 29

get_icmp_check, 29 get_icmp_code, 29 get_icmp_type, 29 Icmp_analysis, 29 icmp_check, 30 icmp_code, 30 icmp_type, 30 Icmp_analysis, 29 icmph, 30

icmp_check

Icmp_analysis, 30 icmp_code

Icmp_analysis, 30 icmp_type

Icmp_analysis, 30 icmph

Icmp_analysis, 30 init_Analysis

simtl_analysis.cpp, 57 simtl_analysis.h, 59

initial

DB_control, 27 IP_ALEN

simtl_analysis.h, 59 Ip_analysis, 31

get_ip_check, 32 get_ip_dest, 32 get_ip_fragoff, 32 get_ip_hlen, 32 get_ip_id, 32 get_ip_proto, 32 get_ip_source, 32 get_ip_tos, 32 get_ip_totlen, 33 get_ip_ttl, 33 get_ip_ver, 33 Ip_analysis, 32 ip_check, 33 ip_destination, 33 ip_fragoff, 33 ip_hlen, 33 ip_id, 33 ip_proto, 33 ip_source, 33 ip_tos, 33 ip_totlen, 34 ip_ttl, 34 ip_ver, 34 Ip_analysis, 32 iph, 34 p_point, 34

ip_check

Ip_analysis, 33 ip_destination

Ip_analysis, 33 ip_fragoff

Ip_analysis, 33 ip_hlen

Ip_analysis, 33 ip_id

Ip_analysis, 33 ip_proto

Ip_analysis, 33 ip_source

Ip_analysis, 33 ip_tos

Ip_analysis, 33 ip_totlen

Ip_analysis, 34 ip_ttl

Ip_analysis, 34 ip_ver

Ip_analysis, 34 iph

Ip_analysis, 34 itemModel

DB_control, 28 itemSelected

Simtl_dir, 39

label

Ui_MainWindow, 50 label_2

Ui_MainWindow, 50 layoutWidget

Ui_MainWindow, 51 layoutWidget1

Ui_MainWindow, 51 layoutWidget2

Ui_MainWindow, 51

main

main.cpp, 54 main.cpp

main, 54 makeDir

Simtl_dir, 39 menubar

Ui_MainWindow, 51 moc_simtl_dir.cpp

qt_meta_data_Simtl_dir, 55

qt_meta_stringdata_Simtl_dir, 55 moc_simtl_gui.cpp

qt_meta_data_Simtl_gui, 56

qt_meta_stringdata_Simtl_gui, 56

okDir

Simtl_dir, 39 operator=

AnalyzerInterface, 20

p_option simtl_analysis.cpp, 57 simtl_analysis.h, 59

p_point Ip_analysis, 34 Packet_analysis, 35 Tcp_analysis, 45

Packet_analysis, 35 data_analysis, 35 i, 35 p_point, 35

packetBox

Ui_MainWindow, 51 packetcnt

Pcap_option, 38 Pcap_option, 36

cflag, 38 chcnt, 38 device, 38 filter_rule, 38 get_cflag, 36 get_chcnt, 36 get_device, 36 get_packetcnt, 37 get_pcap_t, 37 get_rule, 37 packetcnt, 38 Pcap_option, 36 Pcap_option, 36 pd, 38 plus_chcnt, 37 set_cflag, 37 set_chcnt, 37 set_device, 37 set_packetcnt, 37 set_pcap_t, 37 set_rule, 38

pd

Pcap_option, 38 plus_chcnt

Pcap_option, 37

qt_meta_data_Simtl_dir

moc_simtl_dir.cpp, 55 qt_meta_data_Simtl_gui

moc_simtl_gui.cpp, 56 qt_meta_stringdata_Simtl_dir

moc_simtl_dir.cpp, 55 qt_meta_stringdata_Simtl_gui

moc_simtl_gui.cpp, 56 qtable

DB_control, 28

removeDir

Simtl_dir, 40 renameDir

Simtl_dir, 40 retranslateUi

Ui_MainWindow, 49 row_plus

DB_control, 27 rows

DB_control, 28

saveButton

Ui_MainWindow, 51 set_cflag

Pcap_option, 37 set_chcnt

Pcap_option, 37 set_device

Pcap_option, 37 set_itemModel

DB_control, 27 set_model

DB_control, 28 set_options

Simtl_gui, 42 set_packetcnt

Pcap_option, 37 set_pcap_t

Pcap_option, 37 set_rule

Pcap_option, 38 set_table

DB_control, 28 setupUi

Ui_MainWindow, 49 simtl_analysis.cpp

init_Analysis, 57 p_option, 57 uint32_t, 57 simtl_analysis.h

init_Analysis, 59 IP_ALEN, 59 p_option, 59

simtl_db.cpp

db_control, 60 simtl_db.h

db_control, 61 Simtl_dir, 39

changeDir, 39 directory, 40 dirlist, 40 filename, 40 itemSelected, 39 makeDir, 39 okDir, 39 removeDir, 40 renameDir, 40 Simtl_dir, 39 Simtl_dir, 39 simtl_dir

Simtl_gui, 42 Simtl_gui, 41

button_save, 41 button_start, 41 get_table, 41 set_options, 42 simtl_dir, 42 Simtl_gui, 41 Simtl_gui, 41 table_clicked, 42 table_resize, 42 ui, 42

simtl_time.cpp

time_control, 66 simtl_time.h

time_control, 67 start_clock

Time_control, 47 start_time

Time_control, 47 startButton

Ui_MainWindow, 51 statusbar

Ui_MainWindow, 51 stopButton

Ui_MainWindow, 51 sum

DB_control, 28 sumView

Ui_MainWindow, 51

table_clicked

Simtl_gui, 42 table_resize

Simtl_gui, 42 tcp_ack

Tcp_analysis, 45 Tcp_analysis, 43

get_tcp_ack, 44 get_tcp_anumber, 44 get_tcp_check, 44 get_tcp_doff, 44 get_tcp_dport, 44 get_tcp_fin, 44 get_tcp_psh, 44 get_tcp_rst, 44 get_tcp_snumber, 44 get_tcp_sport, 44 get_tcp_syn, 45 get_tcp_upoint, 45 get_tcp_urg, 45 get_tcp_wins, 45 p_point, 45 tcp_ack, 45 Tcp_analysis, 44 tcp_anumber, 45 tcp_check, 45 tcp_doff, 45 tcp_dport, 45 tcp_fin, 45 tcp_psh, 46 tcp_rst, 46 tcp_snumber, 46 tcp_sport, 46 tcp_syn, 46 tcp_upoint, 46 tcp_urg, 46 tcp_wins, 46 Tcp_analysis, 44 tcph, 46 tcp_anumber

Tcp_analysis, 45 tcp_check

Tcp_analysis, 45 tcp_doff

Tcp_analysis, 45 tcp_dport

Tcp_analysis, 45 tcp_fin

Tcp_analysis, 45 tcp_psh

Tcp_analysis, 46 tcp_rst

Tcp_analysis, 46 tcp_snumber

Tcp_analysis, 46 tcp_sport

Tcp_analysis, 46 tcp_syn

Tcp_analysis, 46 tcp_upoint

Tcp_analysis, 46 tcp_urg

Tcp_analysis, 46 tcp_wins

Tcp_analysis, 46 tcph

Tcp_analysis, 46 Time_control, 47 cur_time, 47 get_time, 47 start_clock, 47 start_time, 47 Time_control, 47 Time_control, 47

time_control simtl_time.cpp, 66

simtl_time.h, 67 toolBar

Ui_MainWindow, 51

Ui, 17 ui

Simtl_gui, 42 Ui::MainWindow, 48 Ui_MainWindow, 49

cancelButton, 50 centralwidget, 50 detailBrowser, 50 filterEdit, 50 hboxLayout, 50 hboxLayout1, 50 hboxLayout2, 50 hboxLayout3, 50 label, 50 label_2, 50 layoutWidget, 51 layoutWidget1, 51 layoutWidget2, 51 menubar, 51 packetBox, 51 retranslateUi, 49 saveButton, 51 setupUi, 49 startButton, 51 statusbar, 51 stopButton, 51 sumView, 51 toolBar, 51 widget, 52

uint32_t simtl_analysis.cpp, 57

widget Ui_MainWindow, 52